OpenStudy (anonymous):

State whether every point on the graph is a solution of the equation. Explain why or why not. x – 4y = –4 A. Every point shown on the graph is not a solution of the equation because when the coordinates of the points are substituted into the equation, the equation is not true for all the points. B. Every point shown on the graph is a solution of the equation because when the coordinates of the points are substituted into the equation, the equation is true for all the points.

OpenStudy (anonymous):

pleas help

OpenStudy (aripotta):

can we see the graph?

OpenStudy (aripotta):

or tell us the points on the graph

OpenStudy (anonymous):

-4,0 3,2

OpenStudy (aripotta):

so there's (-4, 0) (0, 1) and (3, 2) our equation is x – 4y = –4 so for each point, we're going to substitute the x and y values into the equation, and see if it's true. i'll do the first point: x – 4y = –4 -4 - 4(0) = -4 -4 - 0 = -4 -4 = -4 so the first point is on the line. can you do the other points?

OpenStudy (anonymous):

thanks